Use the fact that the trigonometric functions are periodic to find the exact value of each expression. Do not use a calculator.
cos 33pi\4.

Answer by Alan3354(69443)

cos 33pi\4.
-------------
32pi/4 = 8pi which is 4 revolutions. Subtract that.
--> cos(pi/4) = sqrt(2)/2
I didn't use a calculator, I know that one.
It's 1/8 of a revolution, or 45º.
If you plot it on a unit circle, x and y are equal, and r is x*sqrt(2)
cos = x/r = sqrt(2)/2